OpenStudy (anonymous):

A line k passes through the points (4,2) and (-4,4).. A second line h passes through the points (-8,1) and (8,-3). Is line k parallel to line h? Why or why not?

OpenStudy (anonymous):

if they have the same slope, then they are parallel

OpenStudy (anonymous):

so since (4-2)/(-4-4)=-1/4 and (-3-1)/(-1+8) = -1/4 they have the same slope, so they are parallel

OpenStudy (anonymous):

Thank y ou

OpenStudy (anonymous):

welcome